Amy travels 450 miles at a certain speed. If the car had gone 15mph faster, the trip would have taken one hour less. Find Amy's speed.
A. 78 mph
B. 75 mph
C. 68 mph
D. 72 mph

Answer by venugopalramana(3286)

LET AMY'S SPEED BE X MPH….TIME TO TRAVEL 450 MILES=450/X HRS.
IF CAR GOES 15 MPH FASTER SPEED = X+15…..TIME TO TRAVEL=450/(X+15)
THIS TIME IS 1 HOUR LESS.SO…450/X-450/(X+15)=1…MULTIPLYING BY LCM X(X+15)
450{X+15-X)=X(X+15)=450*15=6750=X^2+15X
X^2+15X-6750=0=X^2+90X-75X-6750=X(X+90)-75(X+90)=(X-75)(X+90).HENCE X=75….B IS THE ANSWER.
INCIDENTALLY I GOT THROUGH YAHOO 360 A MESSAGE REQUESTING HELP TO YOU FROM YOUR FATHER .I REPLIED IN AFFIRMATIVE TO YAHOO 360 AS I COULD NOT REPLY DIRECTLY O YOUR FATHER A I DONT HAVE HE MAIL ADDRESS.IF YOU GIVE THE ADDRES I CAN CLARIFY BYOUR OTHER DOUBTS TOO.